OpenStudy (anonymous):

Roger jumps an average of 62 times a day. If he jumped 73 times Monday, 65 times Tuesday, and 53 times Thursday, how many times did Roger jump on Wednesday?

OpenStudy (anonymous):

(Monday + Tuesday + Wednesday + Thursday)/4 = 62 (73 + 65 + Wednesday + 53)/4 = 62 (191 + Wednesday) = (62*4) 191 + Wednesday = 248 Wednesday = 248 - 191 Wednesday = 57 jumps

OpenStudy (anonymous):

thaank you so freaking much! =D